# On the microlocal phase space concentration of Schrödinger evolutions

### v1

@@ 4.1. The $\mathcal{A}$-Wigner operator $K_{\mathcal{A}}$ and the Weyl quantization
Metaplectic Wigner distributions enjoy intertwining relations with Weyl operators with symbols in $S^{0}_{0,0}(\mathbb{R}^{2n})$, see [cordero2024wigner, Theorem 5.4]. To be definite, for $a\in\mathcal{S}^{\prime}(\mathbb{R}^{2n})$ we consider the symbols
@@
We will use the following boundedness result for pseudodifferential operators [Toft:2004aa].
@@ 4.2. The $\mathcal{A}$-Wigner wave front set
In [cordero2022wigner], the authors define the so-called $\tau$-Wigner wave front sets by removing the smoothing window in Definition 1.1, with $W_{\tau}$ replacing the STFT. The lack of the mitigating effect of $g$ forced the authors to consider $f\in L^{2}(\mathbb{R}^{n})$. The $\mathcal{A}$-Wigner wave front set is defined as in (1.5) for functions in $L^{2}(\mathbb{R}^{n})$. Here, we define the $\mathcal{A}$-Wigner wave front set and the $\mathcal{A}$-Wigner wave front set of order $s\geq 0$ as follows.
@@
The definition of the $\tau$-Wigner wave front set is a rephrase of the integral definition given in [cordero2022wigner, Definition 5.3] for $\mathcal{A}=\mathcal{A}_{\tau}$, see (47).
Now, as stated in Theorem 5.4 of [cordero2022wigner], given $f\in L^{2}(\mathbb{R}^{n})$ and $\hat{\mathcal{A}}\in\operatorname*{Mp}(2n,\mathbb{R})$
@@ 4.3. $\mathcal{A}$-Wigner microlocality for pseudodifferential operators
This subsection is devoted to proving the so-called microlocality of pseudodifferential operators, in terms of the $\mathcal{A}$-Wigner wave front set, that is Theorem 1.6. For the purpose, we first need the equivalent of [cordero2023wigner, Lemma 5.3] for $\mathcal{A}$-Wigner kernels. Its proof requires the following version of Calderón–Vaillancourt Theorem [CalderonVaillancourt].
@@
We follow the same pattern of the proof of [cordero2023wigner, Lemma 5.3]. Let
@@
(see e.g. [rodino_book, Proposition 0.3]) yields to
@@
We use the same technique that proves [cordero2022wigner, Theorem 1.12]. Let $X_{0}\notin WF_{\mathcal{A}}(f)$. We have to prove that $X_{0}\notin WF_{\mathcal{A}}(\mathrm{Op}^{\mathrm{w}}(a))$. Precisely, if $X_{0}\notin WF_{\mathcal{A}}(f)$ there exists an open conic neighborhood $\Gamma_{X_{0}}$ of $X_{0}$ such that
@@ 5. Wigner microlocality for FIOs with quadratic phase
The aim of this section is to prove Theorem 1.7, which gives the $\mathcal{A}$-Wigner microlocality property (13) for quadratic FIO [cordero2023wigner, cordero2025wigner, cordero2014generalized], defined as follows.
@@
For the rest of this section, $\mathcal{K}$ will be a fixed quadratic FIO, associated with $S$ as in (22), with phase $\Phi$ of the form (153) and amplitude $a\in S_{0,0}^{0}(\mathbb{R}^{2n})$. Its Wigner kernel will be denoted by $\tilde{\kappa}$. We also recall the following representation formula (see [cordero2023wigner]).
@@
Let us write $S(Y)=(S_{1}(Y),S_{2}(Y))$. By applying [cordero2020time, Lemma 6.1.4] we have
@@
(160) $\int_{\Gamma_{X_{0}}}\langle X\rangle^{2s}\lvert W(\mathcal{K}f)(X)\rvert^{2}dX\lesssim\underbrace{\int_{\Gamma_{X_{0}}}\langle X\rangle^{2s}\lvert\tilde{\mathcal{K}}(\psi Wf)(X)\rvert^{2}dX}_{=:I_{1}}+\underbrace{\int_{\Gamma_{X_{0}}}\langle X\rangle^{2s}\lvert\tilde{\mathcal{K}}((1-\psi)Wf)(X)\rvert^{2}dX}_{=:I_{2}}.$
Hence, since $\tilde{\kappa}(X,Y)=h(X,S(Y))$, where $h$ is the kernel of a pseudodifferential operator with symbol in $S_{0,0}^{0}(\mathbb{R}^{2n})$ (see [cordero2023wigner]) we get
@@
6. Microlocal Analysis of phase space interactions
@@ 6.1. The cross $\mathcal{A}$-Wigner wave front set
Let $a,b\in S_{0,0}^{0}(\mathbb{R}^{2n})$ and consider $f,g\in L^{2}(\mathbb{R}^{2n})$. Then
@@
By [cordero2024wigner, Theorem 5.4 (ii)] and Theorem 2.16 we have
@@ 6.2. Wigner Microlocal Analysis of Schrödinger Interaction
we have that the solution of such a problem at time $t\in\mathbb{R}$ can be expressed as (see [hormandersymplectic])
@@
and the corresponding Cauchy problems (17). Their propagators are generalized metaplectic operators, see [cordero2014generalized, cordero2015integral], i.e., they can be written as
@@
and then we prove Theorem 1.9.
@@
Thus, by Theorem 2.16) we have
@@
More generally, we now prove that the same results hold in the case of the $\mathcal{A}$-Wigner wave front set, if $\mathcal{A}\in\operatorname*{Sp}(2n,\mathbb{R})$ is covariant. In this case, recall by Theorem 2.12, that there exists $\Sigma\in\mathcal{S}^{\prime}(\mathbb{R}^{2n})$ such that
@@
Moreover, if we define $\Sigma_{t}(X):=\Sigma(S_{t}(X))$, we denote by $W_{\mathcal{A}_{t}}=Q_{\Sigma_{t}}$ the corresponding Wigner distribution (cf. [cordero2023wigner]). Our generalization follows basically by the following lemma (see [cordero2024wigner, Lemma 7.1]).
@@
Recall the definition of shift-invertibility in Definition 2.13 and the definition of $E_{\mathcal{A}}$ in (56). By [cordero2023characterization, Proposition 4.4], if $W_{\mathcal{A}}$ is covariant then $W_{\mathcal{A}_{t}}$ originating the wavefront $W_{\mathcal{A}_{t}}$ in Theorem 6.5 is also covariant for every $t\in\mathbb{R}$. Here, we prove that also the shift-invertibility of $W_{\mathcal{A}}$ pass to $\mathcal{C}_{t}$ for every $t$. This simplifies considerably the computation of $W_{\mathcal{C}_{t}}$, as exemplified in Example 6.12 below.
@@
respectively [cordero2025sparse]. Let us consider the Wigner wave front of $u_{1}$ and $u_{2}$, whose evolution is governed by the matrix $\mathcal{C}_{t}$ in Theorem 1.10. Explicitly,
@@
To compute $W_{\mathcal{C}_{t}}$ we used that $\mathcal{C}_{t}$ is shift-invertible for every $t\in\mathbb{R}$ by Theorem 6.8. Then, (181) follows by [cordero2024metaplectic, Corollary 4.4] and [cordero2024unified, Lemma 3.6], where $W_{\mathcal{C}_{t}}$ is computed in terms of the short-time Fourier transform and of the Wigner distribution.
@@ 7. Conclusion: a bridge between Wigner and Gabor wave front set
The theory developed in this work heavily relies on the assumption $f\in L^{2}(\mathbb{R}^{n})$ in Definition 1.5, because of which we have been able to define $WF_{\mathcal{A}}(f)$ for any metaplectic Wigner distribution $W_{\mathcal{A}}$. This is due to the fact that $W_{\mathcal{A}}f\in L^{2}(\mathbb{R}^{2n})$ for every $f\in L^{2}(\mathbb{R}^{n})$, so that in principle it makes sense to consider the integral
(183) $\int_{\Gamma}\langle X\rangle^{2s}|W_{\mathcal{A}}f(X)|^{2}dX$
for every $s\in\mathbb{R}$ and open cone $\Gamma$. On the other hand, the classical microlocal analysis via the Gabor wave front $WF_{G}(f)$ can be formulated for $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$, as it is given in terms of the STFT $V_{g}f$, for a window $g\in\mathcal{S}(\mathbb{R}^{n})\setminus\{0\}$, typically a Gaussian, see Definition 1.1. This definition is well-posed thanks to the smoothing effect of $g$, which guarantees $V_{g}f\in\mathcal{C}^{\infty}(\mathbb{R}^{2n})\cap\mathcal{S}^{\prime}(\mathbb{R}^{2n})$.
In Wigner analysis, we are interested in considering $Wf$, thus avoiding resorting to an additional $g$, both because of the physical meaning of $Wf$ and to get information on the ghost-frequencies, that would be otherwise suppressed by the smoothing effect of the window. However, regardless of ghost frequencies, we may be still interested in obtaining information on the order of growth or decay of $W_{\mathcal{A}}f$ (or on the $L^{2}$ integrability) along cones, also when $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$. Reasonably, we may define $\widetilde{WF_{\mathcal{A}}}(f)$ for $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$ by replacing condition (12) with
(184) $W_{\mathcal{A}}f\ast\Phi\in\bigcap_{s\geq 0}L^{2}_{v_{s}}(\Gamma_{X_{0}}),$
where, for instance, $\Phi(X)=2^{n/2}e^{-2\pi|X|^{2}}=Wg(X)$, being $g(y)=e^{-\pi|y|^{2}}$. Condition (184) is clearly not equivalent to (12) and the question arises of what kind of analysis we would get with such a correction. The answer is easy to give. In the case of the classical Wigner distribution, i.e., $\widetilde{WF_{\mathcal{A}}}=\widetilde{WF_{\mathrm{w}}}$, the Gaussian filtering $Wf\ast\Phi$ yields the so-called Husimi distribution [Husimi]:
@@
thereby retrieving $\widetilde{WF_{\mathrm{w}}}=WF_{G}$. In conclusion, the Wigner analysis of wave fronts introduced by Cordero and Rodino intersects Gabor wave front sets when extended to tempered distribution by Gaussian smoothing. This strengthens Wigner analysis as a parallel strategy for the study of evolution operators, propagation of singularities and energy concentration. By working on $Wf$, the gap left by the absence of a window function is filled by the physical interpretability of the Wigner distribution $Wf$ as a quasi-probability distribution, together with its ability to reveal finer microlocal features otherwise suppressed by windowed transforms. At the same time, we do not move far from the classical framework: the familiar Gabor setting and its classical results are recovered through a simple Gaussian convolution.
@@ Acknowledgements
The authors thank prof. Luigi Rodino for reading the manuscript and providing significant suggestions and improvements.
The authors are affiliated to the Gruppo Nazionale per Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA). Gianluca Giacchi is also supported by the SNSF starting grant “Multiresolution methods for unstructured data” (TMSGI2 211684).

### v2

@@ 4.1. The $\mathcal{A}$-Wigner operator $K_{\mathcal{A}}$ and the Weyl quantization
Metaplectic Wigner distributions enjoy intertwining relations with Weyl operators with symbols in $S^{0}_{0,0}(\mathbb{R}^{2n})$, see [20, Theorem 5.4]. To be definite, for $a\in\mathcal{S}^{\prime}(\mathbb{R}^{2n})$ we consider the symbols
@@
We will use the following boundedness result for pseudodifferential operators [61].
@@ 4.2. The $\mathcal{A}$-Wigner wave front set
In [29], the authors define the so-called $\tau$-Wigner wave front sets by removing the smoothing window in Definition 1.1, with $W_{\tau}$ replacing the STFT. The lack of the mitigating effect of $g$ forced the authors to consider $f\in L^{2}(\mathbb{R}^{n})$. The $\mathcal{A}$-Wigner wave front set is defined as in (1.5) for functions in $L^{2}(\mathbb{R}^{n})$. Here, we define the $\mathcal{A}$-Wigner wave front set and the $\mathcal{A}$-Wigner wave front set of order $s\geq 0$ as follows.
@@
The definition of the $\tau$-Wigner wave front set is a rephrase of the integral definition given in [29, Definition 5.3] for $\mathcal{A}=\mathcal{A}_{\tau}$, see (47).
Now, as stated in Theorem 5.4 of [29], given $f\in L^{2}(\mathbb{R}^{n})$ and $\hat{\mathcal{A}}\in\operatorname*{Mp}(2n,\mathbb{R})$
@@ 4.3. $\mathcal{A}$-Wigner microlocality for pseudodifferential operators
This subsection is devoted to proving the so-called microlocality of pseudodifferential operators, in terms of the $\mathcal{A}$-Wigner wave front set, that is Theorem 1.6. For the purpose, we first need the equivalent of [19, Lemma 5.3] for $\mathcal{A}$-Wigner kernels. Its proof requires the following version of Calderón–Vaillancourt Theorem [10].
@@
We follow the same pattern of the proof of [19, Lemma 5.3]. Let
@@
(see e.g. [8, Proposition 0.3]) yields to
@@
We use the same technique that proves [29, Theorem 1.12]. Let $X_{0}\notin WF_{\mathcal{A}}(f)$. We have to prove that $X_{0}\notin WF_{\mathcal{A}}(\mathrm{Op}^{\mathrm{w}}(a))$. Precisely, if $X_{0}\notin WF_{\mathcal{A}}(f)$ there exists an open conic neighborhood $\Gamma_{X_{0}}$ of $X_{0}$ such that
@@ 5. Wigner microlocality for FIOs with quadratic phase
The aim of this section is to prove Theorem 1.7, which gives the $\mathcal{A}$-Wigner microlocality property (13) for quadratic FIO [19, 21, 23], defined as follows.
@@
For the rest of this section, $\mathcal{K}$ will be a fixed quadratic FIO, associated with $S$ as in (22), with phase $\Phi$ of the form (154) and amplitude $a\in S_{0,0}^{0}(\mathbb{R}^{2n})$. Its Wigner kernel will be denoted by $\tilde{\kappa}$. We also recall the following representation formula (see [19]).
@@
Let us write $S(Y)=(S_{1}(Y),S_{2}(Y))$. By applying [28, Lemma 6.1.4] we have
@@
(161) $\int_{\Gamma_{X_{0}}}\langle X\rangle^{2s}\lvert W(\mathcal{K}f)(X)\rvert^{2}dX$ $\lesssim\underbrace{\int_{\Gamma_{X_{0}}}\langle X\rangle^{2s}\lvert\tilde{\mathcal{K}}(\psi Wf)(X)\rvert^{2}dX}_{=:I_{1}}$
(162) $+\underbrace{\int_{\Gamma_{X_{0}}}\langle X\rangle^{2s}\lvert\tilde{\mathcal{K}}((1-\psi)Wf)(X)\rvert^{2}dX}_{=:I_{2}}.$
Hence, since $\tilde{\kappa}(X,Y)=h(X,S(Y))$, where $h$ is the kernel of a pseudodifferential operator with symbol in $S_{0,0}^{0}(\mathbb{R}^{2n})$ (see [19]) we get
@@
6. Cross Wigner microlocal analysis
@@ 6.1. The cross $\mathcal{A}$-Wigner wave front set
Let $a,b\in S_{0,0}^{0}(\mathbb{R}^{2n})$ and consider $f,g\in L^{2}(\mathbb{R}^{n})$. Then
@@
By [20, Theorem 5.4 (ii)] and Theorem 2.16 we have
@@ 6.2. Wigner Microlocal Analysis of Schrödinger Interaction
we have that the solution of such a problem at time $t\in\mathbb{R}$ can be expressed as (see [45])
@@
and the corresponding Cauchy problems (17). Their propagators are generalized metaplectic operators, see [23, 26], i.e., they can be written as
@@
We are ready to prove Theorem 1.9.
@@
Thus, by Theorem 2.16 we have
@@
More generally, we now prove that the same results hold in the case of the $\mathcal{A}$-Wigner wave front set, if $W_{\mathcal{A}}$ is covariant. In this case, recall by Theorem 2.12, that there exists $\Sigma\in\mathcal{S}^{\prime}(\mathbb{R}^{2n})$ such that
@@
Moreover, if we define $\Sigma_{t}(X):=\Sigma(S_{t}(X))$, we denote by $W_{\mathcal{A}_{t}}=Q_{\Sigma_{t}}$ the corresponding Wigner distribution (cf. [19]). Our generalization follows basically by the following lemma (see [20, Lemma 7.1]).
@@
Recall the definition of shift-invertibility in Definition 2.13 and the definition of $E_{\mathcal{A}}$ in (56). By [30, Proposition 4.4], if $W_{\mathcal{A}}$ is covariant then $W_{\mathcal{A}_{t}}$ originating the wavefront $W_{\mathcal{A}_{t}}$ in Theorem 6.5 is also covariant for every $t\in\mathbb{R}$. Here, we prove that also the shift-invertibility of $W_{\mathcal{A}}$ pass to $\mathcal{C}_{t}$ for every $t$. This simplifies considerably the computation of $W_{\mathcal{C}_{t}}$, as exemplified in Example 6.12 below.
@@
respectively [18]. Let us consider the Wigner wave front of $u_{1}$ and $u_{2}$, whose evolution is governed by the matrix $\mathcal{C}_{t}$ in Theorem 1.10. Explicitly,
@@
To compute $W_{\mathcal{C}_{t}}$ we used that $\mathcal{C}_{t}$ is shift-invertible for every $t\in\mathbb{R}$ by Theorem 6.8. Then, (183) follows by [16, Corollary 4.4] and [22, Lemma 3.6], where $W_{\mathcal{C}_{t}}$ is computed in terms of the short-time Fourier transform and of the Wigner distribution.
@@ 7. Conclusion: a bridge between Wigner and Gabor wave front set
The theory developed in this work heavily relies on the assumption $f\in L^{2}(\mathbb{R}^{n})$ in Definition 1.5, because of which we have been able to define $WF_{\mathcal{A}}(f)$ for any metaplectic Wigner distribution $W_{\mathcal{A}}$. This is due to the fact that $W_{\mathcal{A}}f\in\mathcal{C}(\mathbb{R}^{2n})$ for every $f\in L^{2}(\mathbb{R}^{n})$, and it makes sense to consider the integrals
(185) $\int_{\Gamma}\langle X\rangle^{ps}|W_{\mathcal{A}}f(X)|^{p}dX$
for every $p>0$, $s\in\mathbb{R}$ and open cone $\Gamma$. On the other hand, the classical microlocal analysis via the Gabor wave front $WF_{G}(f)$ can be formulated for $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$, as it is given in terms of the STFT $V_{g}f$, for a window $g\in\mathcal{S}(\mathbb{R}^{n})\setminus\{0\}$, typically a Gaussian, see Definition 1.1. This definition is well-posed thanks to the smoothing effect of $g$, which guarantees $V_{g}f\in\mathcal{C}(\mathbb{R}^{2n})\cap\mathcal{S}^{\prime}(\mathbb{R}^{2n})$.
In Wigner analysis, we are interested in considering $Wf$, thus avoiding resorting to an additional $g$, both because of the physical meaning of $Wf$ and to get information on the ghost-frequencies, that would be otherwise suppressed by the smoothing effect of the window. However, regardless of ghost frequencies, we may be still interested in obtaining information on the order of growth or decay of $W_{\mathcal{A}}f$ (or on the $L^{2}$ integrability) along cones, also when $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$. Reasonably, for fixed $p>0$, we may define $\widetilde{WF_{\mathcal{A}}}(f)$ for $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$ by replacing condition (12) with
(186) $W_{\mathcal{A}}f\ast\Phi\in\bigcap_{s\geq 0}L^{p}_{v_{s}}(\Gamma_{X_{0}}),$
where, for instance, $\Phi(X)=2^{n/2}e^{-2\pi|X|^{2}}=Wg(X)$, being $g(y)=e^{-\pi|y|^{2}}$. For $p=2$, condition (186) is clearly not equivalent to (12) and the question arises of what kind of analysis we would get with such a correction. The answer is easy to give. In the case of the classical Wigner distribution, i.e., $\widetilde{WF_{\mathcal{A}}}=\widetilde{WF_{\mathrm{w}}}$, the Gaussian filtering $Wf\ast\Phi$ yields the so-called Husimi distribution [48]:
@@
thereby retrieving $\widetilde{WF_{\mathrm{w}}}=WF_{G}$, for $p=1$. In conclusion, the Wigner analysis of wave fronts introduced by Cordero and Rodino intersects Gabor wave front sets when extended to tempered distribution by Gaussian smoothing. This strengthens Wigner analysis as a parallel strategy for the study of evolution operators, propagation of singularities and energy concentration. By working on $Wf$, the gap left by the absence of a window function is filled by the physical interpretability of the Wigner distribution $Wf$ as a quasi-probability distribution, together with its ability to reveal finer microlocal features otherwise suppressed by windowed transforms. At the same time, we do not move far from the classical framework: the familiar Gabor setting and its classical results are recovered through a simple Gaussian convolution.
@@ Acknowledgements
The authors thank prof. Luigi Rodino for reading the manuscript and providing significant suggestions and improvements. The authors are affiliated to the Gruppo Nazionale per Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA). Gianluca Giacchi is also supported by the SNSF starting grant “Multiresolution methods for unstructured data” (TMSGI2 211684).